# Autonomous Stabilization of Floquet States Using Static Dissipation

### v2

@@ I Introduction
However, Floquet engineering is always limited by heating. There are three conceptually distinct mechanisms for heating. Consider a single degree of freedom with a few energy levels subject to Floquet driving. The stationary states of the system (known as quasienergy states) involve admixtures of both ground and excited states. Accounting for dissipation and dephasing in the laboratory frame leads to both excitation and de-excitation processes in the quasienergy state basis, and thus hot steady states in this basis. Second, restricting to a few levels is always an approximation in a physical system. Higher order processes often allow the system to absorb energy from the drive and leak out of the intended operating subspace. Third, with many interacting degrees of freedom, the system can absorb energy from the drive and heat up, even if the first two mechanisms are suppressed Lazarides et al. (2014); Ponte et al. (2015); Mori et al. (2018).
@@
One route to stabilizing more general, but still time-independent, states is to engineer new dissipation terms distinct from the native dissipators Poyatos et al. (1996); Harrington et al. (2022). For example, in superconducting qubit architectures, two photon loss has been used to stabilize bosonic Kerr-Cat qubits Leghtas et al. (2015); Puri et al. (2017); Grimm et al. (2020), while selective photon addition can stabilize code spaces in a Cat code Gertler et al. (2021). Collective dissipators can also stabilize target entangled states of atoms Plenio et al. (1999); Kraus et al. (2008); Verstraete et al. (2009); Barreiro et al. (2011); Lin et al. (2013); Shankar et al. (2013); Brown et al. (2022). In principle, periodic modulation of engineered dissipators could implement spontaneous emission into a specific quasienergy state. Such an approach is however highly fine-tuned, requiring knowledge of the target quasienergy state at all times during the drive cycle.
@@
A different route to stabilization is to actively extract heat from the system continuously Home et al. (2009a, b); Raghunandan et al. (2020); Petiziol and Eckardt (2022); Wang and Safavi-Naeini (2023); Murch et al. (2012); Lu et al. (2017); Ma et al. (2019); Li et al. (2024) or stroboscopically Lloyd (1996); Sarovar and Milburn (2005); Mi et al. (2024) into an auxiliary system. This type of approach has been used to stabilize arbitrary static single-qubit states Murch et al. (2012); Lu et al. (2017); Ma et al. (2019) and entangled states of two qubits Li et al. (2024). These methods have also been extended to Floquet systems in the limit of high-frequency driving, where they stabilize the ground state space of static effective Hamiltonians. Refs. Petiziol and Eckardt (2022); Wang and Safavi-Naeini (2023) numerically demonstrated such stabilization, while Ref. Mi et al. (2024) experimentally cooled into the ground state of an effective many-qubit Hamiltonian by cycling between a Trotterized Hamiltonian evolution step and an auxiliary qubit reset step.
@@
Figure 1: Static dissipation stabilizes Floquet states. (a) Schematic of the augmented system consisting of a spin with intrinsic dissipation rate $\Gamma$ (dotted wavy arrow) subject to a strong drive of strength $B_{0}$ (orange), coupled to an auxiliary cavity with coupling strength $g$ (light blue) and photonic decay rate $\kappa$ (solid wavy arrow). (b) Quasienergy level structure of the augmented system. The representative quasienergy states, |ϕ±ketsubscriptitalic-$\ket{\phi_{\pm}}$, are shown for $n=0,1$ photons in the cavity. Far from resonance, photon decay drives the system to $n=0$ while preserving the spin state. On resonance however, the cavity coupling $g$ hybridizes spin states with different photon numbers, allowing the state |$\ket{\phi_{+}}$ in the $n=0$ sector to convert to |$\ket{\phi_{-}}$ and decay. When the spin decay rate $\Gamma$ is much weaker, the unique steady state remains close to |$\ket{\phi_{-},n=0}$.
@@
Here we show that, regardless of drive-frequency regime, a static coupling to an unmodulated but lossy auxiliary system can cool a Floquet system into a target quasienergy state. This method is autonomous, requiring no active feedback, and exploits no microscopic knowledge of the time-dependence of the target quasienergy state.
@@
The remainder of the manuscript is organized as follows. In Sec. II, we describe the physical setup involving a periodically modulated spin coupled to a lossy cavity, which showcases the autonomous stabilization effect. In Sec. III, we present experimental evidence for stabilization in this physical setup in the low-frequency (adiabatic) regime. We turn to the general theory in Sec. IV, and provide numerical evidence for stabilization in all frequency regimes. Lastly, in Sec. V, we show that dissipative stabilization enhances the performance of the topological photon pump Nathan et al. (2019); Long et al. (2022).
@@ II The Physical Set-up
This hierarchy is realized when the driving is strong, and the cavity-QED coupling is in the intermediate to strong coupling regime. The representative quasienergies in Eq. (4) depend on the drive frequency regime.
@@ III Experimental Observation of Stabilization in the Adiabatic Regime
Figure 2: Dissipative stabilization of a quasienergy state in the adiabatic limit. The panels show the probability of being in the $e$ or $\ket{\uparrow}$ state (obtained from fast ramp-out measurements) after evolution under the Hamiltonian $H_{B}^{+}$ with $B_{0}=2\pi\times 80$ MHz and $\omega_{\mathrm{mod}}=2\pi\times 0.75$ MHz for three different cavity detunings. The orange (blue) curve is obtained by starting with the qubit aligned (anti-aligned) with the field, which we denote as |$\ket{\phi_{+}}_{0}$ and |$\ket{\phi_{-}}_{0}$, respectively. When $g=\kappa=\Gamma=0$, the curves are sinusoidal with unit amplitude and are exactly out-of-phase. (a) Far from resonance ($\delta=6.15g=2\pi\times 80$ MHz), the curves oscillate out-of-phase for all times, but with a decaying amplitude. At long times, both states are close to a 50-50 mixture in the $z$-basis. Dashed line indicates a fit to the |$\ket{\phi_{+}}_{0}$ data with extracted time constant $T_{d}=10.2\ \mu$s. (b) At intermediate detuning $\delta=2.31g=2\pi\times 30$ MHz, the blue trace shows persistent oscillations, while the orange trace first decays to zero, and re-emerges as a curve with the same amplitude and phase as the blue curve. As the steady state contrast is not unity, the stabilized state is a mixture with greater weight on the $|\phi_{-}\rangle$ state. (c) Closer to resonance, $\delta=0.77g=2\pi\times 10$ MHz, the blue curve is the attractor with a high contrast in the steady state, indicating that $|\phi_{-}\rangle$ is stabilized with high fidelity. The steady-state oscillations persist for far longer than any of the system coherence times, for example, the qubit $T_{1}$ decay shown for reference in black. For details on error bar computation, see Appendix  H.
Figure 3: Onset of stabilization in the quasienergy basis in the adiabatic limit. Measurements of the |$\ket{\phi_{-}}$ population (averaged over two periods) as a function of evolution time. Diamond (circle) markers denote |$\ket{\phi_{+}}$ (|$\ket{\phi_{-}}$) initial state preparation for four different detuning values from the cavity. Curves are labeled by the detuning $\delta/2\pi=(\Delta-B_{0})/2\pi$. Side panels show the applied $2\pi\times 80$ MHz rotating magnetic field (solid line) and the $\delta=0$ ring for each detuning (dashed line). The shaded ring indicates the region of ±$\pm g=\pm 2\pi\times 13$ MHz around cavity resonance where the qubit states can hybridize strongly with the cavity. As $\delta\rightarrow 0$, the system steady state approaches the pure state |$\ket{\phi_{-}}$ with increasing fidelity. Exponential fits to the data are shown in dashed (dotted) lines for preparation in |$\ket{\phi_{-}}$ (|$\ket{\phi_{+}}$).
@@
In this section, we present an experimental implementation of the model in Eq. (3) in the adiabatic limit and demonstrate stabilization of a quasienergy state using a transmon qubit Blais et al. (2021); Krantz et al. (2019) coupled to a superconducting cavity. We show that, when the resonance condition in Fig. 1(b) holds, the transmon is stabilized to a chosen instantaneous eigenstate for timescales that far exceed the native $T_{1}$ and $T_{2}$ decoherence times of the qubit.
@@
In the adiabatic regime, the resonance condition is alternatively visualized in the $B_{x}-B_{z}$ plane. The trajectory traced out by $\vec{B}(t)$ on the plane is a closed curve (black curves, right panels of Fig. 3), while the resonance condition $\delta=\Delta-|\vec{B}(t)|=0$ defines a circle of radius $\Delta$ (dotted, right panels of Fig. 3). Dissipative stabilization occurs in the vicinity of every intersection point of the two curves. In the simple model in Eq. (1), stabilization thus occurs at all times during the protocol, provided $\Delta\approx B_{0}$. However, this is unnecessary; as long as the trajectory traced out by $\vec{B}(t)$ spends a significant fraction of time within roughly $g$ of the $\delta=0$ ring, photonic dissipation stabilizes the chosen quasienergy state. The stabilized cavity-state boosting protocol presented in Sec. V uses this property in an essential manner.
@@
The dynamics of the augmented qubit-cavity system are most transparent if the initial state of the spin coincides with a quasienergy state. In the adiabatic regime, this is an instantaneous eigenstate of $H_{B}$ at the start of the protocol. Since generating $H_{B}$ requires strong driving fields, pulsed preparation of the instantaneous eigenstates is challenging, and we instead make use of an adiabatic preparation scheme. An optional $\pi$-pulse initializes the qubit in the bare |$\ket{g}$ or |$\ket{e}$ states, which are then mapped to |$\ket{\phi_{+}}$ and |$\ket{\phi_{-}}$, respectively, by the adiabatic preparation step. After evolution under $H_{B}^{+}$ for a variable amount of time, the final state is mapped back to the bare transmon basis using either a fast ramp out protocol where the fields are shut off as quickly as possible to preserve the instantaneous qubit populations in the $\sigma_{z}$ basis, or by a slow ramp out protocol where the field is rotated to a fixed point on the Bloch sphere to map adiabatic states to the bare qubit states before measurement. See Appendix C for details of the full waveforms and the different ramp out protocols.
Both fast ramp-out measurements, shown in Fig. 2 and slow ramp-out measurements, shown in Fig. 3 display clear signatures of stabilization to the |$\ket{\phi_{-}}$ state when $\delta\approx 0$. When $\delta\gg g$, the differential loss rate between |$\ket{\phi_{+},0}$ and |$\ket{\phi_{-},1}$ is negligible and the dynamics tends to a (slighly biased) mixed state. However, when $\delta\lesssim g$, significant steady-state coherence in the |$\ket{\phi_{-}}$ state is observed, regardless of whether the initial state is prepared in |$\ket{\phi_{+}}$ or |$\ket{\phi_{-}}$.
@@
Fast ramp-out measurements, such as those shown in Fig. 2, approximately measure the population of the state obtained at the end of the evolution under $H_{B}^{+}$ in the $\sigma_{z}$-basis on the qubit. The two instantaneous eigenstates, |$\ket{\phi_{+}}$ and |$\ket{\phi_{-}}$, manifest as ±$\pm\sin(\omega_{\mathrm{mod}}t)$ oscillations in the probability of qubit excitation after a hold time $t$. In the far detuned case, the cavity induces negligible preferential loss, and the two states evolve with a $\pi$ phase shift before eventually decaying to a mixed state due to dissipative processes on the qubit, as shown in Fig. 2(a). Near resonance, if the qubit is initially prepared in the |$\ket{\phi_{-}}$ state, it remains close to |$\ket{\phi_{-}}$ indefinitely, leading to persistent oscillations [blue curves in Fig. 2(b-c)]. However, qubits initially prepared in |$\ket{\phi_{+}}$ behave qualitatively differently. The initial oscillation damps out, followed by a high-contrast persistent oscillation with the same phase as the |$\ket{\phi_{-}}$ state [orange curves in Fig. 2(b-c)]. These steady-state oscillations persist for timescales far exceeding both the atomic and photonic decoherence times of the system. The crossover from an out-of-phase oscillation with the |$\ket{\phi_{-}}$ state to a persistent in-phase oscillation is most clearly visible at intermediate detuning, such as that shown in Fig. 2(b) where $\delta\approx 2g$.
The slow ramp-out protocol implements a complementary measurement in the instantaneous eigenbasis of the system. These measurements, shown in Fig. 3 provide a direct measure of the |$\ket{\phi_{-}}$ population at the end of the evolution under $H_{B}^{+}$. Neglecting admixture from the one-photon sector in the cavity and imperfect diabaticity of the ramp out, the value of $\langle\sigma_{z}\rangle$ obtained from fast ramp-out measurements is approximately $P(\phi_{+})\sin(\omega_{\mathrm{mod}}t)-P(\phi_{-})\sin(\omega_{\mathrm{mod}}t)$. The slow ramp out measurements in Fig. 3 are therefore related to the contrast of the oscillations in Fig. 2. Away from resonance, both initial states decay to a mixed state with a small bias, $P(\phi_{-})\approx 1/2$. As $\delta\rightarrow 0$, the steady-state approaches a pure state with $P(\phi_{-})\approx 1$.
@@
Near resonance, the cavity increases loss during the adiabatic preparation phase, and as a result, the states starting from |$\ket{g}$ or |$\ket{e}$ end up with significant weight on |$\ket{\phi_{-}}$, even at $t=0$ (yellow curves in Fig. 3). When this weight is near or above $50\%$, the corresponding fast ramp-out data is not expected to be initially out-of-phase. This is borne out by Fig. 2(c).
Appendix F presents the detuning dependence of the steady-state $P(\phi_{-})$ and the timescale for relaxation to the steady state. There are two notable features. First, in addition to a peak at $\delta=0$, $P(\phi_{-})$ is maximal at $\delta=-B_{0}/2$ (or $\Delta=B_{0}/2$). This is a consequence of stabilization by a two photon resonance. Second, peaks in $P(\phi_{-})$ go hand-in-hand with dips in the timescale for relaxation to the steady state. This is a consequence of the larger photonic loss rate $\kappa$, and not the smaller spin dissipation rate $\Gamma$, controlling the decay to the steady state. Both features re-appear in Sec. IV.

### v3

@@ I Introduction
However, Floquet engineering is always limited by heating, or more correctly entropy production. There are three conceptually distinct mechanisms for heating. Consider a single degree of freedom with a few energy levels subject to Floquet driving. The stationary states of the isolated system (known as quasienergy states) involve admixtures of both ground and excited states. Accounting for dissipation and dephasing in the laboratory frame leads to both excitation and de-excitation processes in the quasienergy state basis, and thus mixed steady states with larger entropy than the stationary states of the isolated system. Second, restricting to a few levels is always an approximation in a physical system. Higher order processes often allow the system to absorb energy from the drive, obtain weight on states outside of the intended operating subspace and thus heat up. Third, with many interacting degrees of freedom, the system can absorb energy from the drive and locally become more mixed, even if the first two mechanisms are suppressed Lazarides et al. (2014); Ponte et al. (2015); Mori et al. (2018).
@@
One route to stabilizing more general, but still time-independent, states is to engineer new dissipation terms distinct from the native dissipators Poyatos et al. (1996); Harrington et al. (2022). For example, in superconducting qubit architectures, two photon loss has been used to stabilize bosonic Kerr-Cat qubits Leghtas et al. (2015); Lescanne et al. (2020); Putterman et al. (2022), while selective photon addition can stabilize code spaces in a Cat code Gertler et al. (2021). Collective dissipators can also stabilize target entangled states of atoms Plenio et al. (1999); Kraus et al. (2008); Verstraete et al. (2009); Barreiro et al. (2011); Lin et al. (2013); Shankar et al. (2013); Brown et al. (2022). In principle, periodic modulation of engineered dissipators could implement spontaneous emission into a specific quasienergy state. Such an approach is however highly fine-tuned, requiring knowledge of the target quasienergy state at all times during the drive cycle.
@@
A different route to stabilization is to actively extract heat from the system continuously Home et al. (2009a, b); Raghunandan et al. (2020); Petiziol and Eckardt (2022); Wang and Safavi-Naeini (2023); Murch et al. (2012); Lu et al. (2017); Ma et al. (2019); Li et al. (2024) or stroboscopically Lloyd (1996); Sarovar and Milburn (2005); Mi et al. (2024) into an auxiliary system. This type of approach has been used to stabilize arbitrary static single-qubit states Murch et al. (2012); Lu et al. (2017); Ma et al. (2019) and entangled states of two qubits Shankar et al. (2013); Li et al. (2024). These methods have also been extended to Floquet systems in the limit of high-frequency driving, where they stabilize the ground state space of static effective Hamiltonians. Refs. Petiziol and Eckardt (2022); Wang and Safavi-Naeini (2023) numerically demonstrated such stabilization, while Ref. Mi et al. (2024) experimentally cooled into the ground state of an effective many-qubit Hamiltonian by cycling between a Trotterized Hamiltonian evolution step and an auxiliary qubit reset step.
@@
Figure 1: Static dissipation stabilizes Floquet states. (a) Schematic of the augmented system consisting of a spin with intrinsic dissipation rate $\Gamma$ (dotted wavy arrow) subject to a strong circularly polarized drive of strength $B_{0}$ (orange), coupled to an auxiliary cavity with coupling strength $g$ (light blue) and photonic decay rate $\kappa$ (solid wavy arrow). (b) Quasienergy level structure of the augmented system. The representative quasienergy states, |ϕ±ketsubscriptitalic-$\ket{\phi_{\pm}}$, are shown for $n=0,1$ photons in the cavity. Far from resonance, photon decay drives the system to $n=0$ while preserving the spin state. On resonance however, the cavity coupling $g$ hybridizes spin states with different photon numbers, allowing the state |$\ket{\phi_{+}}$ in the $n=0$ sector to convert to |$\ket{\phi_{-}}$ and decay. When the spin decay rate $\Gamma$ is much weaker than $\kappa$, the unique steady state remains close to |$\ket{\phi_{-},n=0}$.
@@
Here we show that, regardless of drive-frequency regime, a static coupling to an unmodulated but lossy auxiliary system can cool a Floquet system into an isolated target quasienergy state. This method is autonomous, requiring no active feedback, and exploits no microscopic knowledge of the time-dependence of the target quasienergy state. It instead introduces static loss to spectral regions with unwanted states. As a result, the method applies to a broad class of time-dependent Hamiltonians, well beyond the particular examples considered in this article, as long as the target state is not in a dense level spectrum. In particular, the method does not rely either on exact resonance of quasienergies of the non-target states, nor on any resonances in the instantaneous spectrum.
@@
The remainder of the manuscript is organized as follows. In Sec. II, we describe the physical setup involving a periodically modulated spin coupled to a lossy cavity, which showcases the autonomous stabilization effect. In Sec. III, we present experimental evidence for stabilization in this physical setup in the low-frequency (adiabatic) regime. We turn to the general theory in Sec. IV, and provide numerical evidence for stabilization in all frequency regimes. Lastly, in Sec. V, we show that dissipative stabilization enhances the performance of the topological photon pump Nathan et al. (2019); Long et al. (2022). We conclude in Sec. VI, where we discuss further applications beyond two-level systems.
@@ II The Physical Set-up
To effectively cool the system to a pure state, the cooling rate of the system $\kappa$ must exceed the other decay rates in the system ($\Gamma$) which can act as heating rates in the quasi-energy basis. While a large $\kappa$ is desirable for faster cooling, exceeding the qubit-cavity coupling strength ($\kappa>g$) leads to a decoupling of the cavity from the main system, analogous to the bad cavity limit in cavity QED systems or the quantum Zeno effect Misra and Sudarshan (1977); Itano et al. (1990). While the drive frequency $\omega_{\mathrm{mod}}$ does not enter the hierarchy in Eq. (4) explicitly, the quasienergies ϵ±subscriptitalic-$\epsilon_{\pm}$ depend on $\omega_{\mathrm{mod}}$. Thus, the scale of $\omega_{\mathrm{mod}}$ is indirectly constrained in terms of the other parameters by requiring that the quasienergy difference $|\epsilon_{+}-\epsilon_{-}|$ be large. The hierarchy in Eq. (4) is realized when the driving is strong, and the cavity-QED coupling is in the intermediate to strong coupling regime. The representative quasienergies in Eq. (4) depend on the drive frequency regime.
@@ III Experimental Observation of Stabilization in the Adiabatic Regime
In this section, we present an experimental implementation of the model in Eq. (3) in the adiabatic limit (where $\omega_{\mathrm{mod}}\ll B_{0}$) and demonstrate stabilization of a quasienergy state using a transmon qubit Blais et al. (2021); Krantz et al. (2019) coupled to a superconducting cavity. We show that, when the resonance condition in Fig. 1(b) holds, the transmon is stabilized to a chosen instantaneous eigenstate for timescales that far exceed the native $T_{1}$ and $T_{2}$ decoherence times of the qubit.
Figure 2: Onset of stabilization in the quasienergy basis in the adiabatic limit. Measurements of the |$\ket{\phi_{-}}$ population (averaged over two periods) as a function of evolution time. Diamond (circle) markers denote |$\ket{\phi_{+}}$ (|$\ket{\phi_{-}}$) initial state preparation for four different detuning values from the cavity. The error bars are smaller than the markers (see appendices B and H for measurement details). Curves are labeled by the detuning $\delta/2\pi=(\Delta-B_{0})/2\pi$. Side panels show the applied $2\pi\times 80$ MHz rotating magnetic field (solid line) and the $\delta=0$ ring for each detuning (dashed line). The shaded ring indicates the region of ±$\pm g=\pm 2\pi\times 13$ MHz around cavity resonance where the qubit states can hybridize strongly with the cavity. As $\delta\rightarrow 0$, the system steady state approaches the pure state |$\ket{\phi_{-}}$ with increasing fidelity. Exponential fits to the data are shown in dashed (dotted) lines for preparation in |$\ket{\phi_{-}}$ (|$\ket{\phi_{+}}$), with decay times ranging from 12 $\mu$s in the far-detuned regime, to 3 $\mu$s on resonance, consistent with the expected resonant decay rate of roughly $2/\kappa=3.8\,\mu$s. The full detuning dependence of the stabilization time can be found in Appendix F.
Figure 3: Dissipative stabilization of a quasienergy state in the adiabatic limit. The panels show the probability of being in the $e$ or $\ket{\uparrow}$ state (obtained from fast ramp-out measurements, see Appendix C for details) after evolution under the Hamiltonian $H_{B}^{+}$ with $B_{0}=2\pi\times 80$ MHz and $\omega_{\mathrm{mod}}=2\pi\times 0.75$ MHz for three different cavity detunings. The orange (blue) curve is obtained by starting with the qubit aligned (anti-aligned) with the field, which we denote as |$\ket{\phi_{+}}_{0}$ and |$\ket{\phi_{-}}_{0}$, respectively. When there is perfect adiabatic following (and no photon or qubit decay), these two curves are sinusoidal with unit amplitude and are exactly out-of-phase. (a) Far from resonance ($\delta=6.15g=2\pi\times 80$ MHz), the curves oscillate out-of-phase for all times, but with a decaying amplitude. At long times, both states are close to a 50-50 mixture in the $z$-basis. Dashed line indicates a fit to the |$\ket{\phi_{+}}_{0}$ data with extracted time constant $T_{d}=10.2\ \mu$s. (b) At intermediate detuning $\delta=2.31g=2\pi\times 30$ MHz, the blue trace shows persistent oscillations, while the orange trace first decays to zero, and re-emerges as a curve with the same amplitude and phase as the blue curve. As the steady state contrast is not unity, the stabilized state is a mixture with greater weight on the $|\phi_{-}\rangle$ state. (c) Closer to resonance, $\delta=0.77g=2\pi\times 10$ MHz, the blue curve is the attractor with a high contrast in the steady state, indicating that $|\phi_{-}\rangle$ is stabilized with high fidelity. The steady-state oscillations persist for far longer than any of the system coherence times, for example, the qubit $T_{1}$ decay shown for reference in black. For details on error bar computation, see Appendix H.
@@
In the adiabatic regime, the resonance condition is alternatively visualized in the $B_{x}-B_{z}$ plane. The trajectory traced out by $\vec{B}(t)$ on the plane is a closed curve (black curves, right panels of Fig. 2), while the resonance condition $\delta=\Delta-|\vec{B}(t)|=0$ defines a circle of radius $\Delta$ (dotted, right panels of Fig. 2). Dissipative stabilization occurs in the vicinity of every intersection point of the two curves. In the simple model in Eq. (1), stabilization thus occurs at all times during the protocol, provided $\Delta\approx B_{0}$. However, this is unnecessary; as long as the trajectory traced out by $\vec{B}(t)$ spends a significant fraction of time within roughly $g$ of the $\delta=0$ ring, photonic dissipation stabilizes the chosen quasienergy state, as discussed in Appendix E. The stabilized cavity-state boosting protocol presented in Sec. V uses this property in an essential manner.
@@
The dynamics of the augmented qubit-cavity system are most transparent if the initial state of the spin coincides with a quasienergy state. In the adiabatic regime, this is an instantaneous eigenstate of $H_{B}$ at the start of the protocol. Since generating $H_{B}$ requires strong driving fields, pulsed preparation of the instantaneous eigenstates is challenging, and we instead make use of an adiabatic preparation scheme. An optional $\pi$-pulse initializes the qubit in the bare |$\ket{g}$ or |$\ket{e}$ states, which are then mapped to |$\ket{\phi_{+}}$ and |$\ket{\phi_{-}}$, respectively, by the adiabatic preparation step. After evolution under $H_{B}^{+}$ for a variable amount of time, the final state is mapped back to the bare transmon basis using either (i) a slow ramp-out protocol in which the field is rotated to a fixed point on the Bloch sphere to map the adiabatic states to the bare qubit states before measurement, or (ii) a fast ramp-out protocol in which the fields are shut off as quickly as possible to preserve the instantaneous qubit populations in the $\sigma_{z}$ basis. See Appendix C for details of the full waveforms and the different ramp out protocols.
The slow ramp-out measurements, such as those shown in Fig. 2, approximately measure the population in the instantaneous eigenbasis of the system: |$\ket{\phi_{-}},\ket{\phi_{+}}$. In this basis, the measurement axis rotates along with the applied field, removing the time dependence of the rotating field. Far from resonance, both initial states (|$\ket{\phi_{-}}_{0}$ and |$\ket{\phi_{+}}_{0}$) undergo decay to a mixed state with $P(\phi_{-})\approx 0.5$, albeit with a small bias due to imperfections in the adiabatic ramp out. Strikingly, the rate at which the system decoheres is not limited by the native $T_{2}$ phase coherence time (approximately $0.5\,\mu$s) but instead is much closer to the $T_{1}$ time of $13\,\mu$s.
Near resonance ($|\delta|\leq g$), the states |$\ket{\phi_{+},0}$ and |$\ket{\phi_{-},1}$ hybridize leading to preferential loss for |$\ket{\phi_{+}}$ states through the cavity. This loss asymmetry shifts the steady-state of the system from a mixed state to an increasingly pure state, |$\ket{\phi_{-},0}$, as the detuning approaches zero [yellow curves in Fig. 2]. In this regime, if the qubit is initialized in |$\ket{\phi_{-}}_{0}$, it will stay in that state indefinitely whereas a qubit initialized in |$\ket{\phi_{+}}_{0}$ will be dissipatively driven to the |$\ket{\phi_{-}}$ state. In particular, the rate at which the system reaches steady state increases as the detuning approaches zero leading to a faster and stronger stabilization effect on resonance.
As the detuning increases, the steady state fidelity $P(\phi_{-})$ decreases, while the steady state relaxation time increases (see Appendix F for full detuning dependence). The inverse correlation between the steady state fidelity and relaxation time arises from the trade-off between predominantly cavity induced loss (at a rate $\kappa$) near resonance and intrinsic qubit dissipation ( at a rate $\Gamma$) far off resonance. Interestingly, we observe another peak (dip) in the steady state fidelity (relaxation time) at $\delta=-B_{0}/2$ corresponding to a two photon resonance. This qualitative behavior remains even outside the adiabatic limit probed in the experiment, as shown in the following section.
The fast-ramp out measurements, shown in Fig. 3, perform a measurement in the bare basis of the qubit: $\langle\sigma_{z}\rangle$. In this basis, the instantaneous eigenstates of $H_{B}^{+}$ acquire a periodic time-dependence as the applied field rotates in the $xz$-plane of the Bloch sphere. Neglecting admixture from the one-photon sector in the cavity and imperfect diabaticity of the ramp out, the value of $\langle\sigma_{z}\rangle$ obtained from the fast ramp-out measurement is approximately $P(\phi_{+})\sin(\omega_{\mathrm{mod}}t)-P(\phi_{-})\sin(\omega_{\mathrm{mod}}t)$. Far from resonance, states initially prepared in |$\ket{\phi_{-}}_{0}$ versus |$\ket{\phi_{+}}_{0}$ therefore evolve with a $\pi$-phase shift, as shown in Fig. 3(a). The oscillatory response damps out as the both states decay to a mixed state. Just as with the slow ramp-out data, the decay time is significantly longer than the native phase-coherence times, indicating a reduced sensitivity to frequency noise on the qubit transition.
@@
Near resonance, if the qubit is initialized in |$\ket{\phi_{-}}_{0}$, it will remain close to |$\ket{\phi_{-}}$ indefinitely, leading to persistent oscillations [blue curves in Fig. 3(b-c)]. However, qubits initially prepared in |$\ket{\phi_{+}}_{0}$ behave qualitatively differently. The initial oscillation damps out, followed by a high-contrast persistent oscillation with the same phase as the |$\ket{\phi_{-}}$ state [orange curves in Fig. 3(b-c)]. These steady-state oscillations persist for timescales far exceeding both the qubit and photonic decoherence times of the system. The crossover from an out-of-phase oscillations to a persistent in-phase oscillation is most clearly visible at intermediate detuning, such as that shown in Fig. 3(b), where $\delta\approx 2g$.
@@
Due to increased loss from the cavity near resonance during the adiabatic state preparation phase, the states starting from both |$\ket{g}$ or |$\ket{e}$ end up with significant weight on |$\ket{\phi_{-}}$, even at $t=0$ (yellow curves in Fig. 2). When this weight is near or above $50\%$, the corresponding fast ramp-out data is not expected to display initial out-of-phase oscillations, as seen in Fig. 3(c).